OpenStudy (anonymous):

What is the equation of a line that passes through (-3, -5) and has a slope of -2?

OpenStudy (anonymous):

Equation of the line: y=mx + b, where y = your y coordinate, m = your slope and x = your x coordinate.

OpenStudy (anonymous):

so they are asking for the equation of a line, which is in the form of y= mx + b Since we are given the slope, which is m=-2, then the only thing we have to find is b. for that we do the following, we know that (-3,-5), is a (x,y) point on the line so: y=-2x +b; -5=-2(-3) + b -5=6+b b=-11 so the equation of the line is: y=-2x-11

OpenStudy (anonymous):

Therefore, -5 = -2(-3) + b b = -5 - 6 = -11 Eqn: Y = -2x - 11